\begin{thebibliography}{00}

\bibitem{AGV}
Arnold, V.I., Gusein-Zade, S.M., Varchenko, A.N., Singularities
  of Differentiable Maps. vols 1/2, Monographs in Math., \textbf{ 82/83},
Birkh\"auser, Basel (1985/1988)

\bibitem{Art}  E. Artal, J. Carmona, J. I. Cogolludo, M. Escario, J. Fernández de Bobadilla, I. Luengo, A. Melle, Topological and arithmetical properties of rational plane curves. Mathematical contributions in honor of Professor Enrique Outerelo Domínguez (Spanish), 223–242, Homen. Univ. Complut., Editorial Complutense, Madrid, 2004

\bibitem{BCG} I. Bauer, F. Catanese, F. Grunewald, Chebycheff and Belyi polynomials, dessins d'enfants, Beauville surfaces and group theory. Mediterr. J. Math. 3 (2006), no. 2, 121--146.

\bibitem{Chm} S.V.Chmutov, Examples of projective surfaces with many singularities. J.Algebraic Geom. 1  (1992), 191--196.


\bibitem{CD} A.~D.~R.~Choudary, A.~Dimca,  Koszul complexes and hypersurface singularities, 
        Proc. Amer. Math. Soc. 121 (1994), 1009-1016. 



\bibitem{Cog} J. I. Cogolludo-Agustín,  Topological invariants of the complement to arrangements of rational plane curves, Mem. Amer. Math. Soc. 159 (2002), no. 756.




\bibitem{D1} A. Dimca, {\em Singularities and Topology of Hypersurfaces}, 
Universitext, Springer-Verlag, 1992.


\bibitem{DS} A. Dimca, M. Saito, A generalization of Griffiths' theorem on rational integrals, Duke Math. J. 135(2006),303-326.

\bibitem{DSW} A. Dimca,  M. Saito, L. Wotzlaw, A generalization of Griffiths' theorem on rational integrals II,  Michigan Math. J. 58(2009),  603--625.

\bibitem{DSt} A. Dimca, G. Sticlaru, Koszul complexes and pole order filtrations,  arXiv:1108.3976. 



\bibitem{GP} G.-M. Greuel, G. Pfister,  A Singular Introduction to Commutative Algebra (with contributions by O. Bachmann, C. Lossen, and H. Schönemann). Springer-Verlag 2002 (second edition 2007). 

\bibitem{Gr}
Ph.~Griffiths, On the period of certain rational integrals I,
II, Ann. Math. 90 (1969), 460--541.

\bibitem{GH} Ph. Griffith, J. Harris, Principles of Algebraic Geometry.
Wiley, New York (1978)

\bibitem{Ha} R. Hartshorne: Algebraic Geometry, GTM 52, Springer 1977.

\bibitem{KPR}  P.-V. Koseleff, D. Pecker, F. Rouillier,
   Computing Chebyshev knots diagrams, arXiv:1001.5192v2.

\bibitem{PS} C. Peters, J. Steenbrink, {\em Mixed Hodge Structures}, 
Ergeb. der Math. und ihrer Grenz. 3. Folge 52,
Springer, 2008.

\bibitem{ST} H. Schenck, S. O. Tohaneanu, Freeness of Conic-Line Arrangements in $\mathbb P^2$,
 Commentarii. Mathematici Helvetici 84 (2009), 235-258.

\bibitem{Xu} Y. Xu, Lagrange interpolation on Chebyshev points of two variables, 
    J. Approx. Theory, 87 (1996), p. 220--238.

\bibitem{Co} CoCoA: a system for doing Computations in Commutative Algebra.
Available at http://cocoa.dima.unige.it

\bibitem{Sing} Decker, W.; Greuel, G.-M.; Pfister, G.; Sch{\"o}nemann, H.: \newblock {\sc Singular} {3-1-3} --- {A} computer algebra system for polynomial computations.\newblock {http://www.singular.uni-kl.de} (2011).



\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amscd,amssymb}
\usepackage[arrow,matrix]{xy}
\usepackage[plainpages,backref,urlcolor=blue]{hyperref}

\theoremstyle{plain}
\newtheorem{thm}[subsection]{Theorem}
\newtheorem{lem}[subsection]{Lemma}
\newtheorem{prop}[subsection]{Proposition}
\newtheorem{cor}[subsection]{Corollary}

\theoremstyle{definition}

\newtheorem{definition}[subsection]{Definition}
\newtheorem{ex}[subsection]{Example}
\newtheorem{question}[subsection]{Question}

\numberwithin{equation}{section}
\newcommand{\F}{{\mathcal F}}

\newcommand{\B}{{\mathcal B}}
\newcommand{\QQ}{{\mathcal Q}}

\newcommand{\CC}{{\mathcal C}}

\newcommand{\E}{{\mathcal E}}

\newcommand{\al}{{\alpha}}
\newcommand{\be}{{\beta}}

\newcommand{\R}{\mathbb{R}}
\newcommand{\C}{\mathbb{C}}

\newcommand{\PP}{\mathbb{P}}

\DeclareMathOperator{\im}{im}

\DeclareMathOperator{\mult}{mult}

\DeclareMathOperator{\codim}{codim}

\begin{document}

\title [Chebyshev curves, free resolutions and rational curve arrangements]
{Chebyshev curves, free resolutions and rational curve arrangements }

\author[Alexandru Dimca]{Alexandru Dimca$^1$}
\address{Institut Universitaire de France et Laboratoire J.A. Dieudonn\'e, UMR du CNRS 6621,
                 Universit\'e de Nice Sophia-Antipolis,
                 Parc Valrose,
                 06108 Nice Cedex 02,
                 France}
\email{dimca@unice.fr}

\author[Gabriel Sticlaru]{Gabriel Sticlaru}
\address{Faculty of Mathematics and Informatics,
Ovidius University,
Bd. Mamaia 124, 900527 Constanta,
Romania}
\email{gabrielsticlaru@yahoo.com }
\thanks{$^1$ Partially supported by the French-Romanian Programme LEA Math-Mode
and  ANR-08-BLAN-0317-02 (SEDIGA)}

\subjclass[2000]{Primary 14H45, 14H50; Secondary 32S35}

\keywords{Chebyshev polynomial, singularities, Milnor algebra, free resolutions, mixed Hodge structure}

\begin{abstract} First we construct a free resolution for the Milnor (or Jacobian) algebra $M(f)$ of
a complex projective Chebyshev plane curve $\CC_d:f=0$ of degree $d$. In particular, this resolution implies that the dimensions of the graded components $M(f)_k$ are constant for $k \geq 2d-3.$

Then we show that the Milnor algebra of a nodal plane curve $C$ has such a behaviour if and only if all
the irreducible components of $C$ are rational.

For the Chebyshev curves, all of these components are in addition smooth, hence they are lines or conics and explicit factorizations are given in this case.

\end{abstract}

\maketitle


\section{Introduction and statement of results} \label{sec:intro}

Let $T_d(x)=\cos (d \arccos (x))$ be the Chebyshev polynomial of degree $d$, some of whose properties will be recalled in Section 2 below.

The main of these properties is that $T_d(x)$ has only two critical values, namely $\pm 1$.
This was used by S. V. Chmutov to construct complex projective hypersurfaces with a large number of nodes, i.e. $A_1$-singularities, see \cite{AGV}, volume 2, p. 419 and \cite{Chm}. For another occurence of Chebyshev polynomials in modern algebraic geometry see \cite{BCG}.

In this paper we consider the complex Chebyshev curves $\CC_d$ defined in the complex projective plane $\PP^2$ by the equation $T_d(x,y,z)=0$, where $T_d(x,y,z)$ is the polynomial obtained by homogenization of the polynomial $F^+_d(x,y)=T_d(x)+T_d(y)$, hence exactly the 1-dimensional case of Chmutov's hypersurfaces.

It turns out that these curves are unions of lines and conics in general position, see Corollaries \ref{corA} and \ref{corB}. Hence their study may be of interest in the theory of rational curve arrangements, see for instance \cite{Art}, \cite{Cog}, \cite{ST}. There are interesting relations with knot theory as well, see \cite{KPR}.

Let $S=\C[x,y,z]$ be the graded ring of polynomials in $x,y,z$ with complex coefficients and denote by $S_r$ the vector space of homogeneous polynomials in $S$ of degree $r$.
For any polynomial $f \in S_r$ we define the {\it Jacobian ideal} $J_f \subset S$ as the ideal spanned by the partial derivatives $f_x,f_y,f_z$ of $f$ with respect to $x,y,z$ and the graded Milnor (or Jacobian) algebra
\begin{equation} 
\label{eq1}
M(f)=S/J_f.
\end{equation}
The study of such Milnor algebras is related to the singularities of the corresponding projective curve $C_f:f=0$, see \cite{CD}, as well as to the mixed Hodge theory of 
the curve $C_f$ and of its complement $\PP^2 \setminus C_f$, see the foundational article by Griffiths \cite{Gr} and also \cite{DS}, \cite{DSW}, all three papers treating the case of hypersurfaces in $\PP^n$.

The Hilbert-Poincar\'e series of a graded $S$-module $N$ of finite type is defined by 
\begin{equation} 
\label{eq2}
HP(N)(t)= \sum_{k\geq 0} \dim N_kt^k
\end{equation}
and it is known, see for instance \cite{GP}, to be a rational function of the form 

\begin{equation} 
\label{eq3}
HP(N)(t)=\frac{P(N)(t)}{(1-t)^3}.
\end{equation}
In other words, to determine the series $HP(N)(t)$ it is enough to determine the polynomial $P(N)(t)$.

The best way to understand a Milnor algebra $M(f)$ is to construct a free resolution. Our first main result gives such a resolution for the case $f=T_d(x,y,z)$. The result depends on the parity of $d$.

\begin{thm}
\label{thm1}

\noindent (i) If $d=2m$ is even, then there is a free resolution of the Milnor algebra $M=M(T_d(x,y,z))$
of the form
$$ 0 \to S^m(-4m+1) \to S^{m-1}(-4m+3)\oplus S^3(-4m+2)\to S^3(-2m+1) \to S \to M \to 0.$$
In particular,
$$P(M)(t)=1-3t^{2m-1}+(m-1)t^{4m-3}+3t^{4m-2}-mt^{4m-1}.$$

\medskip

\noindent (ii) If $d=2m+1$ is odd, then there is a free resolution of the Milnor algebra $M=M(T_d(x,y,z))$
of the form
$$ 0 \to S^m(-4m-1) \to S^{m}(-4m+1)\oplus S^2(-4m)\to S^3(-2m) \to S \to M \to 0.$$
In particular,
$$P(M)(t)=1-3t^{2m}+mt^{4m-1}+2t^{4m}-mt^{4m+1}.$$

\end{thm}

Here $S^k=S \oplus...\oplus S$, the direct sum of $k$ copies of $S$, is endowed with the usual grading, namely $(S^k)_r=S_r \oplus...\oplus S_r$ and $(S^k(e))_r=(S^k)_{e+r}$. Moreover, linear morphisms $u:N \to N'$ between graded modules are supposed to preserve the grading, i.e. $u(N_r) \subset N'_r$.

Note that in both cases $\deg P(M)(t)=2d-1$ and $P(M)(t)$ is divisible by $(t-1)^2$. To see this, one 
may use Theorem 5.3.7 in \cite{GP}, since the Krull dimension of $M$ is clearly one, as the corresponding zero-set $V(J_f)$ is a union of lines in $\C^3$.
It follows that in fact one may write 
$$HP(M)=\frac{Q(M)(t)}{1-t}$$
where $Q(M)(t)$ is a polynomial of degree $2d-3$.
Using the relation between Hilbert-Poincar\'e series and Tjurina numbers obtained in \cite{CD}, we get the following.

\begin{cor}
\label{cor1}
\noindent (i) If $d=2m$ is even, then the Chebyshev curve $\CC_d$ has $2m(m-1)$ nodes as singularities and
$\dim M_k=2m(m-1)$
for $k \geq 2d-3$.
\medskip

\noindent (ii) If $d=2m+1$ is odd, then the Chebyshev curve $\CC_d$ has $2m^2$ nodes as singularities and
$\dim M_k=2m^2$
for $k \geq 2d-3$.

\end{cor}

The results proved in \cite{CD} for hypersurfaces with isolated singularities imply that we have stabilization at the step $k_0$, i.e. the dimensions $\dim M_k$ become constant for $k \geq k_0$
with $k_0 =3(d-2)+1$ in general, hence a much weaker result than in the case of Chebyshev curves.

It is surprising that the stabilization at $k_0=2d-3$ for a nodal plane curve has a clear-cut geometric description given in the following, which is the second main result of our note. 

\begin{thm}
\label{thm2}
Let $C \subset \PP^2$ be a nodal curve and $f=0$ be a reduced equation for $C$.

Then the following are equivalent.

\noindent (i) The sequence of dimensions $\dim M(f)_k$ stabilizes at $k_0=2d-3$.

\noindent (ii) $\dim M(f)_{2d-3}=n(C)$, the total number of nodes, alias $A_1$-singularities of the curve $C$. 

\noindent (iii) Any irreducible component $C_i$  of $C$ is a rational curve. In particular, if $C_i$ is smooth, then $C_i$ is either a line or a conic. 

\end{thm}

In other words, $(ii)$ is a numerical test to decide whether a given nodal curve $C$ is an arrangement of rational curves, which is the same as asking that $C$ is covered by a finite number of rational parametrizations.

\bigskip

The proof of Theorem \ref{thm1} is elementary, based on a detailed analysis of the geometry of the Chebyshev curve $\CC_d$ and of an associated curve $\B_d$, defined by the difference of two Chebyshev polynomials. On the other hand, the proof of Theorem \ref{thm2} uses in two points subtle facts of mixed Hodge theory, once in the proof of Lemma \ref{lem1} and then at the end of the proof, see the formula \eqref{eqDSW}. The study of the relations between the Hilbert-Poincar\'e series of the Milnor algebra $M(f)$ and Hodge theory is continued in our recent preprint \cite{DSt}.

Numerical experiments with the CoCoA package \cite{Co} and the Singular package \cite{Sing} have played a key role in the completion of this work.

\section{Basic facts on Chebyshev polynomials} \label{sec:two}

The $d$-th Chebyshev polynomial $T_d(x)$ has $d-1$ critical points, namely $\lambda_k=\cos (k\pi/d)$ for $k=1,...,d-1$. Hence $T_d'(\lambda_k)=0$ and $T_d(\lambda_k)=(-1)^k$. Note also that $T_d'' (\lambda_k) \ne 0$ for $d \geq 2$.

It follows that, for $d=2m+1$ odd,  the critical values $\pm 1$ are both attained $m$ times.
When $d=2m$, the critical value $ 1$ is attained $m-1$ times, while $-1$ is attained $m$ times.

The Chebyshev curve $\CC_d$ has as singular points exactly the points $(\lambda_p:\lambda_q:1) \in \PP^2$
such that $p+q$ is odd. The above remarks show that all these singularities are nodes (i.e. singularities of type $A_1$) and their total number is as stated in Corollary \ref{cor1}.

Note that all the singularities of the Chebyshev curve $\CC_d$ are in fact situated in the affine plane $\C^2 \subset \PP^2$ given by $z=1$. We use $x,y$ as coordinates on this affine plane and freely identify a homogeneous polynomial $f \in S_r$ to a polynomial $g(x,y)=f(x,y,1) \in \C[x,y]_r$,
the set of polynomials in $\C[x,y]$ of degree at most $r$.

The singularities of $\CC_d$ corresponds to the set of points in $\C^2$ given by
\begin{equation} 
\label{eq4}
A_d=\{(\lambda_p,\lambda_q)~~:~~0<p<d,~~0<q<d, ~~p+q \text{ odd } \}.
\end{equation}
We consider also the complementary set
\begin{equation} 
\label{eq5}
B_d=\{(\lambda_p,\lambda_q)~~:~~0<p<d,~~0<q<d, ~~p+q \text{ even } \}.
\end{equation}
Note that the affine curve $\B_d: F_d^-(x,y):=T_d(x)-T_d(y)$ has only nodes as singularities and they are located exactly at the points in $B_d$. The polynomial $F_d^-(x,y)$ has a wonderful factorization,
established in \cite{KPR}. 

\begin{prop}
\label{prop1}
The polynomial $F^-_d(x,y)$ is a product of linear and quadratic factors. More precisely, if we set
$$g_k(x,y)= x^2-2\lambda_{2k}xy+y^2-(1-\lambda_{2k}^2),$$
then one has the following.

\noindent (i) If $d=2m$ is even, then the linear factors are $x-y$ and $x+y$, and the irreducible quadratic factors are $g_k(x,y)$ for $k=1,...,m-1$.

\medskip

\noindent (ii) If $d=2m+1$ is odd, then  the only linear factor is $x-y$, and the irreducible quadratic factors are $g_k(x,y)$ for $k=1,...,m$.

\end{prop}

Recall that $T_d(x)$ is an even (resp. odd) polynomial when $d$ is even (resp. odd). Hence, for $d$ odd, we get
$$F_d^+(x,y)=F^-_d(x,-y).$$
Using Proposition \ref{prop1} we get the following.

\begin{cor}
\label{corA}
If $d=2m+1$ is odd, then the Chebyshev polynomial $F_d^+(x,y)$ splits as a product of the linear factor $x+y$ and the quadratic factors $g^+_k(x,y)= x^2+2\lambda_{2k}xy+y^2-(1-\lambda_{2k}^2),$ for $k=1,...,m.$
In particular, the polynomials $F_d^+(x,y)$ and $F^-_d(x,y)$ are affinely equivalent.
\end{cor}

The same method of proof as for Proposition \ref{prop1} yields the following.

\begin{cor}
\label{corB}
If $d=2m$ is even, then the Chebyshev polynomial $F_d^+(x,y)$ splits as a product of the quadratic factors $g^+_k(x,y)= x^2+2\lambda_{2k-1}xy+y^2-(1-\lambda_{2k-1}^2),$ for $k=1,...,m.$
In particular, in this case the polynomials $F_d^+(x,y)$ and $F^-_d(x,y)$ are not affinely equivalent.
\end{cor}

Note that each quadratic form $g_k(x,y)$ defines an ellipse $\E_k$ in the real plane $\R^2$, whose symmetry axes are the lines $L_1:x-y=0$ and $L_2:x+y=0$. In particular, the intersections $L_j \cap \E_k$ consists of two distinct points in $B_d$.

Moreover, two distinct ellipses $\E_k$ and $\E_{\ell}$ intersect in precisely four distinct points given by 
\begin{equation} 
\label{eq6}
(\lambda_{k+\ell}, \lambda_{k-\ell}), ~~ (-\lambda_{k+\ell}, -\lambda_{k-\ell})~~, (\lambda_{k-\ell}, \lambda_{k+\ell}),\text{ and }
(-\lambda_{k-\ell}, -\lambda_{k+\ell}).
\end{equation}
Notice that this $4$ points determine a rectangle, call it $R_{k,\ell}$.

\section{Construction of the free resolutions} \label{sec:3}

Let $\F(B_d)$ be the vector space of $\C$-valued functions defined on $B_d$.

For each positive integer $r$, we have an evaluation map
\begin{equation} 
\label{eq7}
e(d,r): \C[x,y]_r \to \F(B_d)
\end{equation}
obtained by sending a polynomial $h \in \C[x,y]_r$ to the function $x \mapsto h(x)$ for $x \in B_d$.
We have the following interpolation-type result, similar in spirit to \cite{Xu}.

\begin{prop}
\label{prop2}
The map $e(d,r)$ is injective for $r \leq d-3$ and surjective for $r \geq d-2$.

\end{prop}

\proof

To prove the injectivity part, it is enough to treat the case $r=d-3$. Let $Q \in \C[x,y]_{d-3}$ be a nonzero polynomial vanishing at any point in $B_d$ and let $\QQ$ be the corresponding complex plane curve. If $L$ is a line which is an irreducible component of the curve $\B_d$, then $L$ contains exactly $d-1$ points in $B_d$. It follows that
$$\sum_{x \in L\cap \QQ}\mult_x(L, \QQ) \geq d-1>\deg L \cdot \deg Q=d-3.$$
It follows by Bezout's Theorem, see for instance \cite{GH}, p. 172, that $L$ is an irreducible component of $\QQ$.

Consider next the conic (which is the complex version of the ellipse) $\E_k$. It meets any other conic $\E_{\ell}$ for $\ell \ne k$ in four points. 

In the case $d=2m$ even, we get in this way $4(m-2)$ points of this type on $\E_k$, all of them in $B_d$. There are four more points coming from the intersections of $\E_k$ with the lines $L_1$ and $L_2$. In all we get $4(m-1)$
points of intersection between $\E_k$ and $\QQ$. One has
$$\sum_{x \in \E_k \cap \QQ}\mult_x(\E_k, \QQ) \geq 4(m-1)>\deg \E_k \cdot \deg Q=2(2m-3).$$
It follows by Bezout's Theorem that $\E_k$ is an irreducible component of $\QQ$.

In the case $d=2m+1$ even, we get in this way $4(m-1)$ points of intersection with the other conics on $\E_k$. There are $2$ more points coming from the intersections of $\E_k$ with the line $L_1$. In all we get $4m-2$
points of intersection between $\E_k$ and $\QQ$. One has
$$\sum_{x \in \E_k \cap \QQ}\mult_x(\E_k, \QQ) \geq 4m-2>\deg \E_k \cdot \deg Q=2(2m-2).$$
It follows again by Bezout's Theorem that $\E_k$ is an irreducible component of $\QQ$.

The above shows that $Q$ is divisible by the polynomial $F^-_d(x,y)$, which is a contradiction, since
$\deg Q < \deg F^-_d(x,y)$.

\medskip

To prove the surjectivity, it is enough to suppose $r=d-2$. For each point $b \in B_d$ we construct a polynomial $h_b \in \C[x,y]_{d-2}$ such that $h_b(b) \ne 0$ and $h_b(c)=0$ for any $c \in B_d$, $c \ne b$.

Since all the singularities of $\B_d$ are nodes, it follows that at any point $b \in B_d$ there are exactly two irreducible components of $\B_d$ meeting there. It follows that there are three cases to consider.

\medskip

\noindent (i) the point $b$ is the intersection of two lines (then of course $b=(0,0)$, $d=2m$ and the lines are $L_1$ and $L_2$). Define in this case
$$h_b(x,y)=\frac{F^-_d(x,y)}{x^2-y^2}.$$
\medskip

\noindent (ii) the point $b$ is the intersection of one line $L$ (with equation $\ell(x,y)=0$)
and a conic $\E_k$. Let $b'$ be the second intersection point of these two curves and let $\ell'(x,y)=0$ be the equation of a line passing through $b'$ but different from $L$. Then one can take
$$h_b(x,y)=\frac{F^-_d(x,y)\ell'(x,y)}{\ell(x,y) g_k(x,y)}.$$

\medskip

\noindent (iii) the point $b$ is the intersection of two conics $\E_k$ and $\E_{\ell}$. Let $\ell_1(x,y)=0$ and $\ell_2(x,y)=0$ be the equation of the two sides in the rectangle $R_{k, \ell}$
not passing through $b$. Then one can take
$$h_b(x,y)=\frac{F^-_d(x,y)\ell_1(x,y)\ell_2(x,y)}{g_k(x,y)  g_{\ell}(x,y)}.$$

\endproof

Now we prove Theorem \ref{thm1} by constructing rather explicitly the free resolution of the Milnor algebra $M$. We set $f=T_d(x,y,z)$ and use the notations $F_d^-(x,y,z)$, $g_j(x,y,z)$ a.s.o. to denote the homogenized polynomials associated to  $F_d^-(x,y)$, $g_j(x,y)$ a.s.o.

We consider first the case $d=2m$ even. Then we have to construct a resolution
\begin{equation} 
\label{res}
0 \to R_3 \to R_2 \to R_1 \to R_0 \to M \to 0
\end{equation}
where $R_0=S$, $R_1=S^3(-2m+1)$, $R_2=S^{m-1}(-4m+3)\oplus S^3(-4m+2)$ and $R_3= S^m(-4m+1)$.
The first step is the same as in the Koszul complex, namely we define $u_0:R_1 \to R_0$ to be the map
$(a_1,a_2,a_3) \mapsto a_1f_x+a_2f_y+a_3 f_z$.

Then $K_1=\ker u_0$ is the module of linear relations (syzygies) involving the partial derivatives 
$f_x,f_y,f_z$. Among these, there are three trivial relations, coming also from the Koszul complex, namely
$$r_1=(f_y,-f_x,0),~~r_2=(f_z,0,-f_x),~~r_3=(0,f_z,-f_y).$$
For $(c_1,c_2,c_3) \in S^3(-4m+2) $, we define $u_1(c_1,c_2,c_3)=c_1r_1+c_2r_2+c_3r_3$.

Now we look for nontrivial relations, i.e. relations not lying in the submodule $u_1 (S^3(-4m+2))$.
Note that any relation
\begin{equation} 
\label{eq7.5}
a_1f_x+a_2f_y+a_3 f_z=0
\end{equation}
with all $a_j$ homogeneous polynomials in some $S_r$, implies that $a_3$ vanishes on the set $B_d$.
Indeed, $f_x$ and $f_y$ vanishes on the set $A_d \cup B_d$, and $f_z$ is not vanishing at any point in $B_d$.

Let $E(d,r)=\ker e(d,r)$ and recall that it may be thought of as a subspace in $S_r$ by homogenization. Then Proposition \ref{prop2} implies that $E(d,r)=0$ and hence $a_3=0$ if $r <d-2$. But this implies that $a_1=a_2=0$, hence there are no relations in this range.

Consider now the case $r=d-2$. Using Proposition \ref{prop2}, it follows that the vector space $E(d,r)$  has  dimension given by 
$$ \dim \C[x,y]_{d-2} - |B_d|=m(2m-1)-(2m(m-1)+1)=m-1.$$
On the other hand we have obviously the following linear independent elements in $E(d,d-2)$:
\begin{equation} 
\label{eq8}
\al_3^j=\frac{F_d^-(x,y,z)}{g_j(x,y,z)} \text{ for } j=1,...,m-1.
\end{equation}
Since $\al_3^jf_z$ vanishes on $A_d\cup B_d$, using Noether AF+BG Theorem, see for instance \cite{GH}, p. 703, we get the existence
of unique polynomials $\al_1^j$ and $\al_2^j$ in $S_{d-2}$ such that
$$\al_1^jf_x+\al_2^jf_y+\al_3^j f_z=0$$
for $j=1,...,m-1.$
Hence we've got $m-1$ new relations $\rho_j=(\al_1^j,\al_2^j,\al_3^j) \in K_1$ and we use them to define $u_1(b_1,...,b_{m-1})=b_1\rho_1+....+ b_{m-1}\rho_{m-1}$ for any $(b_1,...,b_{m-1}) \in S^{m-1}(-4m+3)$.

In this way the morphism $u_1$ is completely defined and we have to show that any relation in $K_1$ is in fact in the image of $u_1$. So consider a relation as in \eqref{eq7.5}.

\medskip

The first case to check is when $r=d-1$. As above, $a_3 \in E(d,d-1)$, and this vector space has dimension $(m-1)+2m$, 
again by Proposition \ref{prop2}.

Note that $E'=xE(d,d-2)\oplus yE(d,d-2) \oplus zE(d,d-2) \subset E(d,d-1)$ is a $(3m-3)$-dimensional subspace, hence it is enough to find 2 new linearly independent elements.
Recall that $f_x=F_d^-(x,y,z)_x$ and $f_y=-F_d^-(x,y,z)_y$. 

On the other hand we have $F_d^-(x,y,z)=(x-y)(x+y)\prod_{j=1,m-1}g_j(x,y,z)$. It follows that 
$$f_x=F_d^-(x,y,z)_x=F_d^-(x,y,z)( \frac{1}{x-y}+ \frac{1}{x+y} +\sum_{j=1,m-1}\frac{g_j(x,y,z)_x}{g_j(x,y,z)} )$$
and 
$$-f_y=F_d^-(x,y,z)_y=F_d^-(x,y,z)( -\frac{1}{x-y}+ \frac{1}{x+y} +\sum_{j=1,m-1}\frac{g_j(x,y,z)_y}{g_j(x,y,z)} ).$$
It follows that no nontrivial linear combination of $f_x$ and $f_y$ belongs to $E'$, in other words we may use $xg_j$, $yg_j$, $zg_j$, $f_x$ and $f_y$ to get a basis of $E(d,d-1)$.

This shows that starting with any relation in $(K_1)_{d-1}$, we can modify it by a relation in the image of $u_1$ (in fact without using the relation $r_1$) in order to get a relation with $a_3=0$. But then we get $f_x|a_2$ and $f_y|a_1$, i.e. we get a multiple of the relation $r_1$.

\medskip

The next case to investigate is $r=d$.  Again, $a_3 \in E(d,d)$, and this vector space has dimension $(m-1)+2m+2m+1$, 
again by Proposition \ref{prop2}.

We  consider as above the vector space $E''$ spanned by all the product $q\al_3^j$, where $q$ runs through a basis of $S_2$.
But now these products are not linearly independent, in fact by \eqref{eq8}
one clearly has 
$$g_i(x,y,z)\al_3^i=g_j(x,y,z)\al_3^j$$
for any $i,j=1,...,m-1$.
It follows that the relation
$$g_1(x,y,z)\rho_1-g_j(x,y,z)\rho_j$$ 
for $j=2,...,m-1$ has a trivial last component, and as such is a multiple of the relation $r_1$.
It follows that there are homogeneous polynomials $\be_j \in S_1$ such that one has
$$g_1(x,y,z)\rho_1-g_j(x,y,z)\rho_j-\be_jr_1=0,$$
for $j=2,...,m-1$.

Hence we get $m-2$ relations among the fundamental relations $\rho_j$ and $r_k$.
We use them to define the morphism $u_2:R_3\to R_2$ on the first $(m-2)$-components of 
$R_3= S^m(-4m+1)$ by setting 
$$u_2(e_j)=(g_1(x,y,z),..., -g_j(x,y,z),....,-\be_j,0,0)$$
where $e_j$ is the canonical basis of $ S^m(-4m+1)$, $-g_j(x,y,z)$ is placed on the $j$-th component in $S^{m-1}(-4m+3)$ for $j=2,...,m-1$, and $-\be_j$ is placed on the first component of $S^3(-4m+2)$,
and moreover $...$ stand for zero components.
There are two more new relations among the fundamental relations $\rho_j$ and $r_k$.
Using the formulas for $f_x$ and $f_y$ given above, we see that
$$(x+y)(f_x-f_y)=2g_1\al_3^1+ \sum_{j=1,m-1}(x+y)((g_j)_x+(g_j)_y)\al_3^j$$
and
$$(x-y)(f_x+f_y)=2g_1\al_3^1+ \sum_{j=1,m-1}(x-y)((g_j)_x-(g_j)_y)\al_3^j),$$
where we set $g_j=g_j(x,y,z)$, $(g_j)_x=g_j(x,y,z)$ a.s.o. to simplify the notation. 
As above, these relations produce relations among relations:
$$(x+y)(r_2-r_3)+2g_1\rho_1+ \sum_{j=1,m-1} (x+y)((g_j)_x+(g_j)_y)  \rho_j-\gamma_1r_1=0$$
and
$$(x-y)(r_2+r_3)+2g_1\rho_1+ \sum_{j=1,m-1} (x-y)((g_j)_x-(g_j)_y)  \rho_j-\gamma_2r_1=0$$
with $\gamma_k \in S_1$.
Using them, we complete the definition of $u_2$ by setting
$u_2(e_{m-1})$ to be 
$$(2g_1+(x+y)((g_1)_x+(g_1)_y),...., (x+y)((g_{m-1})_x+(g_{m-1})_y)  , -\gamma_1,x+y,-(x+y))$$
and
$$u_2(e_{m})=( 2g_1+(x-y)((g_1)_x-(g_1)_y),...., (x-y)((g_{m-1})_x-(g_{m-1})_y)  ,    -\gamma_2,x-y,x-y).$$
The injectivity of $u_2$ follows from the fact that the corresponding matrix
(regarded as a matrix over the field of rational functions $\C(x,y,z)$) has rank $m$.
Indeed, a nonzero size $m$ minor can be obtained by deleting the first and the $m$-th rows in this
matrix.

Returning to the inclusion $(K_1)_d \subset \im (u_1)_d$, the above discussion shows that $\dim E''=3(m-1)-(m-2)=2m-1$
and that it is enough to add the vector space spanned by $xf_x$ and $yf_x$ to $E''$ to get $E(d,d) $.
We conclude that $(K_1)_d \subset \im (u_1)_d$ as above, by first showing that we may assume $a_3=0$.

We leave to the reader to check in the same way the inclusion $(K_1)_r \subset \im (u_1)_r$, for $r>d$ as well as the inclusion $\ker (u_1) \subset \im(u_2)$.

\bigskip

The proof in the case $d=2m+1$ follows the same pattern, with just one major difference.
To construct the required resolution in this case, namely
\begin{equation} 
\label{r1}
R_*: 0 \to S^m(-4m-1) \to S^{m}(-4m+1)\oplus S^2(-4m)\to S^3(-2m) \to S 
\end{equation}
is the same as constructing first a larger resolution
$$R'_*: 0 \to S^m(-4m-1)\oplus S(-4m) \to S^{m}(-4m+1)\oplus S^3(-4m)\to S^3(-2m) \to S $$
and them simplifying it. 

We construct the larger resolution $R'_*$, because its construction is more natural. To start with, the morphism $u_0:S^3(-2m) \to S $ is the same as above, coming from the Koszul resolution.

Let $r=d-2=2m-1$. Then $E(d,r)$ is an $m$-dimensional vector space with a basis given by $\al_3^j$
defined exactly as above, with $j=1,...,m$. They yield $m$ new relations $\rho_j$ with $j=1,...,m$
which are used, together with the trivial relations $r_k$ introduced above, to define the morphism $u_1:S^{m}(-4m+1)\oplus S^3(-4m)\to S^3(-2m)$.

\medskip

Let $r=d-1=2m$, $E(d,r)$ is a $(3m+1)$-dimensional vector space and the vector subspace $E'$ constructed as above has dimension $3m$.

Note that in this case $F_d^-(x,y,z)=(x-y)\prod_{j=1,m}g_j(x,y,z)$. It follows that 
$$f_x=F_d^-(x,y,z)_x=F_d^-(x,y,z)( \frac{1}{x-y} +\sum_{j=1,m}\frac{g_j(x,y,z)_x}{g_j(x,y,z)} )$$
and 
$$-f_y=F_d^-(x,y,z)_y=F_d^-(x,y,z)( -\frac{1}{x-y} +\sum_{j=1,m}\frac{g_j(x,y,z)_y}{g_j(x,y,z)} ).$$
Hence the {\it key new fact in the case $d$ odd} is the relation
$$f_x-f_y=\sum_{j=1,m}(g_j(x,y,z)_x+g_j(x,y,z)_y)\al_3^j.$$
This gives a relation
$$r_2-r_3+\sum_{j=1,m}(g_j(x,y,z)_x+g_j(x,y,z)_y)\rho_j-\delta_1r_1=0$$
which is used to define a morphism $ S(-4m) \to S^{m}(-4m+1)\oplus S^3(-4m)$ which will describe the action of $u_2$ on the factor $S(-4m)$.

For the remaining part of the morphism $u_2$, namely a morphism $S^m(-4m-1) \to S^{m}(-4m+1)\oplus S^3(-4m)$, we need $m$ relations among the relations $\rho_j, r_k$.
The first $(m-1)$ relations are obtained exactly as above, and they have the form
$$g_1(x,y,z)\rho_1-g_j(x,y,z)\rho_j-\be_jr_1=0,$$
for $j=2,...,m$.
The last relation comes from the equality
$$(x-y)f_x=g_1\al_3^1+\sum_{j=1,m}(g_j)_x\al_3^j$$
and hence has the form
$$(x-y)r_2+g_1\rho_1+\sum_{j=1,m}(g_j)_x\rho_j-\delta_2r_1=0.$$

\medskip

To get the formulas for the Hilbert-Poincar\'e series, we start with the resolution \eqref{res} or \eqref{r1} and get
$$HP(M)(t)=HP(R_0)(t)-HP(R_1)(t)+HP(R_2)(t)-HP(R_3)(t).$$
Then we use the well-known formulas $HP(N\oplus N')(t)=HP(N)(t)+HP( N')(t)$, $HP(N(-r))(t)=t^rHP(N)(t)$
and
$$HP(S)(t)=\frac{1}{(1-t)^3}.$$

\section{On rational curve arrangements} \label{sec:4}

The proof of Theorem \ref{thm2} depends on the following standard general fact.
For general properties of mixed Hodge theory we refer to \cite{PS}, see also the corresponding Appendix in \cite{D1} for a brief introduction.

\begin{prop}
\label{prop3}
Let $C \subset \PP^2$ be a nodal curve and set $U=\PP^2\setminus C$. Let $C=\cup_{j=1,r}C_j$ be the decomposition of $C$ as a union of irreducible components, let $\nu_j:\tilde C_j \to C_j$ be the normalization mappings and set $g_j=g(\tilde C_j)$.
Then one has 
$$\dim Gr _F^1H^2(U,\C)=\sum_{j=1,r}g_j$$
and
$$\dim Gr _F^2H^2(U,\C)=\frac{(d-1)(d-2)}{2}.$$

\end{prop}

The second formula follows also from  \cite{DSW}, Theorem (2.2), but we give here a more elementary
independent proof.

\proof

Assume that the curve $C_j:f_j=0$ has degree $d_j$ and has $n_j$ nodes. Recall the definition of the Hodge-Deligne polynomial of a quasi-projective complex variety $X$
$$P(X)(u,v)=\sum_{p,q}E^{p,q}(X)u^pv^q$$
where $E^{p,q}(X)=\sum_s(-1)^s \dim Gr_F^pGr^W_{p+q}H^s_c(X,\C)$, and the fact that it is additive with respect to constructible partitions, i.e. $P(X)=P(X \setminus Y)+P(Y)$, for a Zariski closed subvariety $Y$ in $X$.

Using the normalization maps $\nu_j$, it follows that
$$P(C_j)=P(C_j \setminus (C_j)_{sing})+ P((C_j)_{sing})=P(\tilde C_j \setminus \{2n_j \text { points }\})+n_j=$$
$$=P(\tilde C_j)-P( \{2n_j \text { points }\})+n_j=uv-g_ju-g_jv+1-n_j.$$
Using the fact that the only intersection points are nodes, we get
$$P(C)=P(C_1 \cup....\cup C_r)=\sum_{j=1,r}P(C_j) - \sum_{1 \leq i<j\leq r} P(C_i \cap C_j)=$$
$$
=ruv-(\sum_{j=1,r}g_j)u-(\sum_{j=1,r}g_j)v+r -(\sum_{j=1,r}n_j)- \sum_{1 \leq i<j\leq r}d_id_j.$$
Next we have again by the additivity
$P(U)=P(\PP^2)-P(C)$ where $P(\PP^2)=u^2v^2+uv +1$.
Now, let's look at the cohomology of the smooth surface $U$. The group $H^4_c(U,\C)$, is dual to the group $H^0(U,\C)$, which is $1$-dimensional of type $(0,0)$. It follows that the contribution of 
$H^4_c(U,\C)$ to $P(U)$ is exactly the term $u^2v^2$.
 
The group $H^3_c(U,\C)$, is dual to the group $H^1(U,\C)$, which is $(r-1)$-dimensional of type $(1,1)$. It follows that the contribution of 
$H^3_c(U,\C)$ to $P(U)$ is exactly the term $-(r-1)uv$.
The remaining terms come from the group $H^2_c(U,\C)$, which is dual to the group $H^2(U,\C)$, and this implies that $H^2(U,\C)$
has only classes of type $(2,1)$, $(1,2)$ and $(2,2)$. The dimension $\dim Gr _F^1H^2(U,\C)$
is the number of independent classes of type $(1,2)$ (which correspond to terms in $u$ in $P(U)$), and this gives the first equality.

To establish the second equality we have to work a little more. The dimension $\dim Gr _F^2H^2(U,\C)$
is the number of independent classes of type $(2,1)$ or $(2,2)$ (which correspond to terms in $v$ or to the constant terms in the polynomial $P(U)$). This yields
\begin{equation} 
\label{future}
\dim Gr _F^2H^2(U,\C)=\sum_{j=1,r}(g_j+n_j-1)+\sum_{1 \leq i<j\leq r}d_id_j+1.
\end{equation}
To complete the proof, recall the formula
$$g_j+n_j=p_a(C_j)=\frac{(d_j-1)(d_j-2)}{2},$$
where $p_a$ denotes the arithmetic genus, see \cite{Ha}, p. 298 and p. 54. The result follows
using the relation $d=\sum_{j=1,r}d_j$.

\endproof

We need another preliminary result. Consider the evaluation map
$$e(C)_k:S_k \to \F(\Sigma(C))$$
where $\Sigma(C) \subset \C^3 \setminus \{0\}$ is a finite set in bijection to the nodes of $C$ under the canonical projection $\C^3 \setminus \{0\} \to \PP^2$, $\F(\Sigma(C))$ is the vector space of complex valued functions defined on $\Sigma(C)$, and $e(C)_k(h)$ is the function $s \mapsto h(s)$ for
$h \in S_k$ and $s \in \Sigma(C)$.
\begin{lem} 
\label{lem1}
The mapping $e(C)_{2d-3}$ is surjective.
\end{lem}

\proof
As in the proof of Proposition \ref{prop2} above, it is enough to construct for each point $a \in \Sigma(C)$ a homogeneous polynomial $h_a \in S_p$ for some $p \leq 2d-3$ such that $h_a(a) \ne 0$
and $h_a(b)=0$ for any $b \in \Sigma(C)$, $b \ne a$.
There are two cases to discuss.

\medskip

(i) $a \in C_i \cap C_j$ for $i \ne j$. Then the intersection $C_i\cap C_j$ consists of exactly $d_id_j$ points and a function $h \in S_q$ with $q=d_i+d_j-1$
is in the ideal $I(C_i \cap C_j)$, i.e. $h$ vanishes on this set $C_i\cap C_j$ if and only if can be written (in an unique way) as a sum $h=h_if_i+h_jf_j$ with $\deg(h_i)=d_j-1$ and 
$\deg(h_j)=d_i-1$. It follows that 
$$\dim (S/I(C_i\cap C_j))_q=d_id_j$$
which shows that the corresponding evaluation map $e_q:S_q \to \F(C_i\cap C_j)$ is surjective.
Hence we can find $h' \in S_q$ such that $h'(a) \ne 0$ and $h'(b)=0$ for all $b \in (C_i\cap C_j)$,
$b \ne a$.
Since $a$ is a smooth point on $C_i$ (resp. $C_j$), there is a partial derivative of $f_i$ (resp. of $f_j$) not vanishing at $a$. Denote these partial derivatives by $g_1$ and $g_2$ respectively.

Then the product
$$h_a=h'g_1g_2 \prod_{k=1,r;~~k\ne i;~~k\ne j}f_k$$
has the required properties and $\deg h_a =d+d_i+d_j -3 \leq 2d-3.$

\medskip

(ii) $a$ is a node on the irreducible curve $C_i$. Using the exact sequences (3.13) on p. 201,
(3.16) on p. 202 (for $n=s=2$) and the Example (3.18) on p. 203 (for $m=1$) in \cite{D1}, it follows that the evaluation map
$$e(C_i):S_{d_i-3} \to \Sigma(C_i)$$
can be identified to the surjective map
$$F^2H^2(U,\C)= P^2H^2(U,\C) \to \oplus_kF^2H^2(U_k \setminus C_i,\C)$$
where $P$ denotes the polar filtration on cohomology and $U_k$ a small disc around each singular point of the curve
$C_i$.

Hence it exists a polynomial $h' \in S_{d_i-3}$ such that  $h'(a) \ne 0$ and $h'(b)=0$ for all other nodes $b$ of the curve $C_i$.
Then the product
$$h_a=h'\prod_{k=1,r;~~k\ne i}f_k$$
has the required properties and $\deg h_a =d-3 \leq 2d-3.$ 
\endproof

Next we recall some notation and a key result from \cite{DSW} in our special case.

Let $I \subset S$ be the graded ideal of polynomials vanishing at all the singular points of $C$.
In fact $I_k$ is exactly the kernel of the evaluation map $e(C)_k$, for all $k \geq 0$.
By definition, note that $J_f \subset I$ and also $\dim (S_k/I_k) \leq n(C)$ for all $k$.
Moreover, we have $\dim (S_k/I_k)=n(C)$ exactly when the corresponding evaluation map
$$e(C)_k:S_k \to \F(\Sigma (C))$$
is surjective. The above Lemma shows that this holds for $k=2d-3$.

It is shown in \cite{DSW}, Theorem (2.2) and subsection (2.3), that for a nodal curve in $\PP^2$
one has a natural isomorphism
\begin{equation} 
\label{eqDSW}
Gr _F^1H^2(U,\C)=(I/J_f)_{2d-3}.
\end{equation}
Moreover, it follows from \cite{CD}, Corollary 8, that there are epimorphisms
\begin{equation} 
\label{eqCD}
M(f)_{q-1} \to M(f)_q
\end{equation}
for $q \geq 2d-3$. In particular, $\codim (J_f)_{2d-3}\geq n(C)$, where $\codim$ refers to codimension with respect to $S_{2d-3}.$
This last fact implies that the claim $(i)$ is equivalent to the claim $(ii)$,
since we know that $M(f)_k=\tau(C)=n(C)$ for $k$ large.

If we assume $(ii)$, then by Lemma \ref{lem1} we have  
$$\codim I_{2d-3} = n(C)= \codim (J_f)_{2d-3},$$
and hence we get $(iii)$.

On the other hand, the condition $(iii)$ is saying that $\codim (J_f)_{2d-3}=\codim I_{2d-3}$
which in view of Lemma \ref{lem1} implies $(ii)$.

\end{document}